OpenStudy (nathaliama13):

The weight of an African elephant is 6 × 10^6 grams; the weight of a tiger is 3 × 10^5 grams. How many times heavier is the elephant than the tiger? 2 grams 20 grams 200 grams 2,000 grams

OpenStudy (nathaliama13):

@answerhelps

OpenStudy (nathaliama13):

@EclipsedStar

OpenStudy (cwrw238):

i guess you must mean 6 * 10^6 for weigh of an elephant !! - otherwise it woudnt be an elephant

OpenStudy (nathaliama13):

I- yes. I changed it. aha

OpenStudy (cwrw238):

lol

OpenStudy (nathaliama13):

Do you have any suggestions on what I should do to solve it?

OpenStudy (cwrw238):

well you need to divide 6*10^6 by 3*10^5

OpenStudy (nathaliama13):

I got it. thank you lots. i'll give you a medal and fan you. ☻

OpenStudy (cwrw238):

if you are not comfortable with scientific notation convert them to a natral number

OpenStudy (cwrw238):

oh ok yw

OpenStudy (nathaliama13):

wait. i'm not quite sure what I got. idk if it's 20 or 200 ;-; i got 2000000000000

OpenStudy (nathaliama13):

@cwrw238

OpenStudy (cwrw238):

6 * 10^6 =6 * 1,000,000 3 * 10^5 = 3 * 100,000

OpenStudy (cwrw238):

so its 6,000,000 / 300,000

OpenStudy (cwrw238):